INVERSION MOTIONS IN THE DIMETHYL ETHER–CARBON DISULFIDE DIMER

\underline{REBECCA A. PEEBLES}, JOSH J. NEWBY, SEAN A. PEEBLES, Department of Chemistry, Eastern Illinois University, 600 Lincoln Ave., Charleston, IL 61920 USA.

The rotational spectrum of the 1:1 dimer formed between dimethyl ether (DME) and CS_{2} has been assigned using Fourier transform microwave spectroscopy. The spectrum shows tunneling splittings of approximately 180 MHz in the c-type transitions and smaller splittings of a few hundred kilohertz in the a-type transitions. The splittings are attributed to a motion of the CS_{2} between the lone pairs of the DME, causing an inversion of the c-component of the dipole moment. The ΔE(0^{−} − 0^{+}) splitting between the two tunneling states is 90.3411(14) MHz. The B and C rotational constants for the two states were fit separately, along with an average A rotational constant, giving values of A = 8099.5906(12) MHz, B_{0} = 604.67510(18) MHz, B_{1} = 604.65215(18) MHz, C_{0} = 582.50781(23) MHz, and C_{1} = 582.55142(23) MHz. A barrier to the tunneling motion has been estimated using both ab initio techniques and the one dimensional model of Meyer.^{a} Dipole moment components of μ_{a} = 0.961(1) MHz, μ_{c} = 1.104(1) MHz, and μ_{total} = 1.464(1) MHz have been measured and are also consistent with the CS_{2} interacting with a DME lone pair.

aR. Meyer, J. Mol. Spectrosc., 76 (1979) 266.